OpenStudy (anonymous):

in derivatives of trigonomic functions ..I need to answer y=tan(4x) ..please help me..

OpenStudy (anonymous):

chain rule... y' = [tan(4x)]' * (4x)' = sec^2(4x) * 4 = 4*sec^2(4x)

OpenStudy (anonymous):

this can also be expressed other ways of course... do you have the answer?